### v1

# Penalty decomposition derivative free method for the minimization of partially separable functions over a convex feasible set

Abstract
In this paper, we consider the problem of minimizing a smooth function, given as finite sum of black-box functions, over a convex set. In order to advantageously exploit the structure of the problem, for instance when the terms of the objective functions are partially separable, noisy, costly or with first-order information partially accessible, we propose a framework where the penalty decomposition approach is combined with a derivative-free line search-based method. Under standard assumptions, we state theoretical results showing that the proposed algorithm is well-defined and globally convergent to stationary points. The results of preliminary numerical experiments, performed on test problems with number of variables up to thousands, show the validity of the proposed method compared with a standard derivative-free line search algorithm. Moreover, it is shown that the method is easily parallelizable and hence capable of taking advantage of parallelization of computation, when possible.
keywords:
Penalty decomposition, Finite-sum, Coordinate partially separable, Derivative-free line search
articletype: ARTICLE
{amscode}
90C56, 90C30, 90C26
1 Introduction
In this work, we are interested in finite-sum optimization problems of the form
$\begin{array}[]{l}\min f(x)=\displaystyle{\sum_{j=1}^{m}}f_{j}(x)\\ \text{s.t. }x\in X\end{array}$ (1)
where $X\subset\mathbb{R}^{n}$ is a closed, convex set and $f_{j}:\mathbb{R}^{n}\to\mathbb{R}$, for $j=1,\ldots,m$ are smooth functions for which, however, we do not have access to first-order information. We also assume that $f$ is bounded below by some value $f^{*}$.
The literature on derivative-free methods is wide and increasing. For insights on general derivative-free approaches we refer the reader to the books [5, 2], the survey [15] and to some seminal papers both for smooth optimization (e.g., [7, 23, 16, 22]) and nonsmooth optimization (e.g., [1, 8, 17]).
For strongly structured problems, like (1), tailored derivative-free methods, have been presented; for instance, in [6] authors exploit the peculiarities of the feasible set (being the convex hull of a given set of atoms) of the problem; partial knowledge of first order information is taken into account in [18]; the partially separable structure of the objective function is addressed in [21].
Drawing inspiration from the latter work, we have considered (1), noting that its structure may deserve attention in designing derivative-free algorithms, more in general, in the following cases:
  1. (i)
    the objective function is coordinate partially separable (CPS), i.e., each function $f_{j}$ only depends on a (relatively small) subset of variables $S_{j}\subseteq\{1,\ldots,n\}$;
  2. (ii)
    the evaluation of some (or all) terms of the objective function may be costly;
  3. (iii)
    the functions $f_{j}$ may be affected by noise differently as result of different simulations or measurements;
  4. (iv)
    the objective function is partially derivative-free, i.e., for some functions $f_{j}$ the gradient is available.
In order to take into account the structure of problem (1) related to the possible cases (i)-(iv), we adopt a penalty decomposition approach that allows to manage cases (i)-(iv) in a unified way.
Briefly, the penalty decomposition approach consists in the introduction of an auxiliary vector of variables - copies in principle of the original ones - and the addition of a penalty term to the objective function, penalizing the distance between each pair of corresponding variables. The introduction of the auxiliary variables allows the decomposition of the problem into smaller independent subproblems; this may be advantageously exploited to take into account the peculiarities of problem (1) in the cases (i)-(iv) discussed above.
In recent years, various structured problems have been handled by a penalty decomposition strategy; for instance, penalty decomposition allows for an efficient management of sparsity constraints [19] and more generally of geometric constraints [13], or to manage the intersection of convex sets [10]. In all of these cases, the availability of first order information was assumed. To our knowledge, derivative-free penalty decomposition methods have not yet been studied, except for the case of [14], where a method for sparse optimization has been designed to minimize a smooth black-box function to a cardinality constraint.
In this work we define a general algorithm model where the penalty decomposition strategy is coupled with a derivative-free line search-based method for the management of subproblems. We state theoretical results - not immediately descending from known results in the literature - for proving global convergence properties of the proposed algorithm and we present its characterization in the specific case of partially separable functions. Finally, we report the results of preliminary computational experiments.
The paper is organized as follows: Section 2 concerns the penalty decomposition reformulation of problem (1), with the convergence analysis of the sequential scheme discussed in Section 2.1; the inner optimization scheme is presented in 3 and formally analyzed in 3.1, whereas the discussion on the stopping condition is provided in 3.2; in Section 4 we address the particular case of CPS problems; numerical experiments are described and discussed in Section 5; we finally give concluding remarks in Section 6.
Notation
We denote by $\Pi_{X}$ the Euclidean projection operator onto the convex set $X$. A point $x\in X$ is stationary for problem (1) if $x=\Pi_{X}[x-\nabla f(x)]$. We denote by $e_{i}$ the $i$-th element of the canonical basis in $\mathbb{R}^{n}$. If $S\subseteq\{1,\ldots,n\}$ is an index set, we denote by $x_{S}$ the subvector of variables of $x$ indexed by $S$.
2 A Penalty decomposition approach
To describe the methodology proposed in this paper and also for subsequently carrying out its convergence analysis, the following problem, that allows for a much simpler notation, can be considered without loss of generality:
$\min_{x}$ $f(x)=f_{1}(x)+f_{2}(x)$ (2)
s.t. $x\in X.$
Following a well-established approach [19, 10, 14, 13], the problem can be reformulated as
$\min_{x,y,z}$ $f_{1}(y)+f_{2}(z)$ (3)
s.t. $x\in X,$
$x-y=0,$
$x-z=0,$
which is clearly equivalent to (2). Problem (3) then allows to introduce the penalty function
$P_{\tau}(x,y,z)=f_{1}(y)+f_{2}(z)+\frac{\tau}{2}\left(\|x-y\|^{2}+\|x-z\|^{2}\right),$
whose (partial) gradients are given by
$\nabla_{x}P_{\tau}(x,y,z)$ $=\tau\left(x-y\right)+\tau\left(x-z\right),$ (4)
$\nabla_{y}P_{\tau}(x,y,z,w)$ $=\nabla_{x}f_{1}(y)-\tau\left(x-y\right),$
$\nabla_{z}P_{\tau}(x,y,z,w)$ $=\nabla_{x}f_{2}(z)-\tau\left(x-z\right).$
A sequential penalty method [3, Ch. 4.2] can then be employed to address problem (3) and thus (2), which iteratively solves (inexactly) subproblems of the form
$\min_{x,y,z\in\mathbb{R}^{n}}$ $P_{\tau_{k}}(x,y,z)$
[…]
for a sequence of values $\{\tau_{k}\}$ such that $\tau_{k}\to\infty$. The subproblem is considered solved if a solution $(x^{k},y^{k},z^{k})$ is found such that
$\|\nabla_{y}P_{\tau_{k}}(x^{k},y^{k},z^{k})\|$ $\leq c_{y}\xi_{k},$ (5)
$\|\nabla_{z}P_{\tau_{k}}(x^{k},y^{k},z^{k})\|$ $\leq c_{z}\xi_{k},$ (6)
$\|x_{k}-\Pi_{X}[x_{k}-\nabla_{x}P_{\tau_{k}}(x^{k},y^{k},z^{k})]\|$ $\leq\xi_{k},$ (7)
for some positive tolerance $\xi_{k}>0$ and constants $c_{y},c_{z}>0$. The resulting algorithm is formally described by Algorithm 1. Note that, formally, the solution $(x^{k-1},y^{k-1},z^{k-1})$ is not used during iteration $k$; however, it is well-known that, with sequential penalty type methods, using the solution of the previous subproblem as a starting point for the next one is often good for efficiency and stability reasons [3, Ch. 4.2]. We now turn to the convergence analysis of the sequential penalty framework.
1 Input: $x^{0},y^{0},z^{0}\in\mathbb{R}^{n},$$\{\xi_{k}\}\subseteq\mathbb{R}_{+},\{\tau_{k}\}\subseteq\mathbb{R}_{+}$, $c_{y},c_{z}>0$
2 for $k=1,2,\ldots$ do
3       Find $(x^{k},y^{k},z^{k})\in X\times\mathbb{R}^{n}\times\mathbb{R}^{n}$ s.t.
$\|\nabla_{y}P_{\tau_{k}}(x^{k},y^{k},z^{k})\|\leq c_{y}\xi_{k},$
$\|\nabla_{z}P_{\tau_{k}}(x^{k},y^{k},z^{k})\|\leq c_{z}\xi_{k},$
$\|x_{k}-\Pi_{X}[x_{k}-\nabla_{x}P_{\tau_{k}}(x^{k},y^{k},z^{k})]\|\leq\xi_{k}.$
Algorithm 1 Sequential Penalty Approach
2.1 Convergence analysis
We start the analysis stating the following reasonable assumption (formally given for the case of $f$ defined as the sum of $m$ terms).
Assumption 1.
Functions $f_{j}:\mathbb{R}^{n}\to\mathbb{R}$, $j=1,\ldots,m$, are coercive.
We now begin stating the formal properties of the algorithm with a preliminary result.
Proposition 2.1.
For any $\tau>0$, the penalty function $P_{\tau}(x,y,z)$ is coercive, i.e., $P_{\tau}(x^{t},y^{t},z^{t})\to+\infty$ if $\|(x^{t},y^{t},z^{t})\|\to\infty$.
Proof.
By Assumption 1, function $f(y,z)=f_{1}(y)+f_{2}(z)$ is also coercive. Assume $\{(x^{t},y^{t},z^{t})\}$ is such that $\|(x^{t},y^{t},z^{t})\|\to\infty$. If either $\|y^{t}\|\to\infty$ or $\|z^{t}\|\to\infty$, by the coercivity of $f(x,y)$, we get that
$\lim_{t\to\infty}P_{\tau}(x^{t},y^{t},z^{t})$ $\geq\lim_{t\to\infty}f(y^{t},z^{t})=+\infty.$
On the other hand, if $\{(y^{t},z^{t})\}$ is bounded, it must be $\|x^{t}\|\to\infty$. Then, we have
$\lim_{t\to\infty}P_{\tau}(x^{t},y^{t},z^{t})$ $\geq\lim_{t\to\infty}f^{*}+\frac{\tau}{2}(\|y^{t}-x^{t}\|^{2}+\|z^{t}-x^{t}\|^{2})=+\infty.$
Corollary 2.2.
Let $\bar{x}\in X$, $\bar{y},\bar{z}\in\mathbb{R}^{n}$, $\tau>0$. The level set
=$\mathcal{L}=\{(x,y,z)\in X\times\mathbb{R}^{n}\times\mathbb{R}^{n}\mid P_{\tau}(x,y,z)\leq P_{\tau}(\bar{x},\bar{y},\bar{z})\}$
is compact.
We now make a second, reasonable assumption on the algorithm.
Assumption 2.
For all $k$, $(x^{k},y^{k},z^{k})$ satisfies
$P_{\tau_{k}}(x^{k},y^{k},z^{k})\leq P_{\tau_{0}}(x^{0},y^{0},z^{0}).$ (8)
Remark 1.
Note that, for any $\tau>0$ and any feasible $(x,y,z)$, we have
$P_{\tau}(x,y,z)=f(x).$
Therefore, if we start with a solution $(x^{0},y^{0},z^{0})$ such that $x^{0}\in X$ and $y^{0}=z^{0}=x^{0}$, condition (8) can certainly be satisfied if at iteration $k$ we generate $(x^{k},y^{k},z^{k})$ by means of a monotone descent method starting at
$\begin{cases}(x^{k-1},y^{k-1},z^{k-1})&\text{if }P_{\tau_{k}}(x^{k-1},y^{k-1},z^{k-1})\leq f(x^{0}),\\ (x^{0},y^{0},z^{0})&\text{otherwise}.\end{cases}$
We are now able to state a first property concerning the sequence $\{(x^{k},y^{k},z^{k})\}$.
Proposition 2.3.
Let $\{(x^{k},y^{k},z^{k})\}$ be a sequence satisfying Assumption 2 . Then $\{(x^{k},y^{k},z^{k})\}$ is a bounded sequence.
Proof.
By Corollary 2.2 we have that the set
$\mathcal{L}_{0}=\{(x,y,z)\in X\times\mathbb{R}^{n}\times\mathbb{R}^{n}\mid P_{\tau_{0}}(x,y,z)\leq P_{\tau_{0}}(x^{0},y^{0},z^{0})\}$
is compact. Assumption 2 implies that
$P_{\tau_{0}}(x^{k},y^{k},z^{k})\leq P_{\tau_{k}}(x^{k},y^{k},z^{k})\leq P_{\tau_{0}}(x^{0},y^{0},z^{0}),$
where the first inequality follows from the fact that $\tau_{0}\leq\tau_{k}$. Therefore, the points of the sequence $\{(x^{k},y^{k},z^{k})\}$ belong the compact set $\mathcal{L}_{0}$ and this concludes the proof. ∎
We finally state the main convergence result for Algorithm 1.
Proposition 2.4.
Let $\{(x^{k},y^{k},z^{k})\}$ be the sequence generated by Algorithm 1 with $\xi_{k}\to 0$ and $\tau_{k}\to\infty$. Then, $\{(x^{k},y^{k},z^{k})\}$ admits limit points and every limit point $(x^{*},y^{*},z^{*})$ is feasible, i.e., $x^{*}=y^{*}=z^{*}$, and $x^{*}$ is a stationary point of problem (2).
Proof.
Consider the vectors $\mu_{y}^{k},\mu_{z}^{k}\in\mathbb{R}^{n}$ defined as follows
$\mu_{y}^{k}=\tau_{k}(x_{k}-y_{k}),\qquad\mu_{z}^{k}=\tau_{k}(x_{k}-z_{k})$ (9)
Recalling (4), we therefore have
$\nabla_{x}P_{\tau_{k}}(x^{k},y^{k},z^{k})$ $=\mu_{y}^{k}+\mu_{z}^{k},$ (10)
$\nabla_{y}P_{\tau_{k}}(x^{k},y^{k},z^{k})$ $=\nabla_{x}f_{1}(y^{k})-\mu_{y}^{k},$
$\nabla_{z}P_{\tau_{k}}(x^{k},y^{k},z^{k})$ $=\nabla_{x}f_{2}(z^{k})-\mu_{z}^{k}.$
Since, by Proposition 2.3, the sequence $\{(x^{k},y^{k},z^{k})\}$ is bounded, it admits limit points. Let $(x^{*},y^{*},z^{*})$ be any limit point, i.e., there exists an infinite subset $K\subseteq\{0,1,\ldots\}$ such that
$\lim_{k\in K,\,k\to\infty}(x^{k},y^{k},z^{k})=(x^{*},y^{*},z^{*})$
Taking the limits in (10) for $k\in K$ and $k\to\infty$, recalling (5)-(6) and that $\xi_{k}\to 0$, we obtain
$\nabla_{x}f_{1}(y^{*})=\mu^{*}_{y}=\lim_{k\to\infty,\,k\in K}\mu_{y}^{k},\qquad\nabla_{x}f_{2}(z^{*})=\mu^{*}_{z}=\lim_{k\to\infty,\,k\in K}\mu_{z}^{k},$ (11)
Furthermore, we can divide both sides of the two equations in (9) by $\tau_{k}$ and take again the limits, to obtain that $x^{*}=y^{*}=z^{*}$. Therefore, we have
$\mu^{*}_{y}+\mu^{*}_{z}=\nabla_{x}f_{1}(y^{*})+\nabla_{x}f_{2}(z^{*})=\nabla f(x^{*}).$ (12)
Using (10), condition (7) can be written as follows
$\|x_{k}-\Pi_{X}[x_{k}-\mu_{y}^{k}-\mu_{z}^{k}]\|\leq\xi_{k}$ (13)
Then, taking the limits in (13) for $k\in K$ and $k\to\infty$, recalling that $\xi_{k}\to 0$, and using (12), we obtain
$x^{*}=\Pi_{X}[x^{*}-\mu_{y}^{*}-\mu_{z}^{*}]=\Pi_{X}[x^{*}-\nabla f(x^{*})].$
Remark 2.
The result in Proposition 2.4 cannot be directly inferred, to the best of our knowledge, by known results in the literature. Indeed, most known results concerning penalty and augmented Lagrangian methods rely on different stationarity conditions for subproblems with lower-level constraints to be handled explicitly out of the penalty function (see, e.g., [4]). On the other hand, the case of projection-based stationarity has been previously considered in [10, 9], but a bond between the sequences $\{\tau_{k}\}$ and $\{\xi_{k}\}$ is required ($\tau_{k}\xi_{k}\to 0$). For this particular setup we thus provide a novel, improved result of convergence that does not need this particular interaction between the two sequences of parameters.
3 A derivative-free alternate minimization algorithm for penalty subproblems
For addressing problems of the form (1), we assumed not to have access to derivatives information. We are thus interested in devising a computationally convenient algorithmic scheme to tackle subproblems (2) in a derivative-free fashion, so that the overall problem can be solved via Algorithm 1. The derivative-free setup poses two main challenges: from the one hand, subproblems have to be tackled without relying on gradient based methods; on the other hand, a maybe more subtle issue lies in the inner stopping condition: we cannot directly measure the size of the gradients $\|\nabla_{y,z}P_{\tau_{k}}(x^{k},y^{k},z^{k})\|$.
We observe that, for a given value of the penalty parameter $\tau$, it then follows that
    fixed $x$, we have two unconstrained separable problems w.r.t. $y$ and $z$;
    fixed $y$ and $z$, we have a problem in $x$ whose globally optimal solution can be determined, being it convex quadratic with a separable objective function subject to convex constraints; the solution is in fact available in closed form in the particular case of bound constraints.
Similarly as in the previous section, we denote by $k$ the counter of the outer iterations of the sequential penalty approach. For a fixed $k$, we can then set up an inner optimization loop where we perform derivative-free inexact optimization w.r.t. $y$ and $z$ and the exact minimization with respect to $x$. We will denote by $\ell$ the iteration counter of the inner cycles. The algorithm is described in Algorithm 2.
1 Input: $\hat{x}^{k-1},\hat{y}^{k-1},\hat{z}^{k-1}\in\mathbb{R}^{n},$ $\tau_{k}>0$
2 Set $\ell=0$
3 $(x^{\ell},y^{\ell},z^{\ell})=(\hat{x}^{k-1},\hat{y}^{k-1},\hat{z}^{k-1})$
4 Set $\tilde{\alpha}^{\ell}_{y}=\tilde{\alpha}^{\ell}_{z}=(1,\ldots,1)\in\mathbb{R}^{n}$
5while stopping condition not satisfied do
6       $(y^{\ell+1},\tilde{\alpha}^{\ell+1}_{y})=$ DF-Search($P_{\tau_{k}}(y,x)$, $x^{\ell}$, $y^{\ell}$, $\tilde{\alpha}^{\ell}_{y}$)
7       $(z^{\ell+1},\tilde{\alpha}^{\ell+1}_{z})=$ DF-Search($P_{\tau_{k}}(z,x)$, $x^{\ell}$, $z^{\ell}$, $\tilde{\alpha}^{\ell}_{z}$)
8       Set $x^{\ell+1}=\arg\min_{x\in X}P_{\tau_{k}}(x,y^{\ell+1},z^{\ell+1})$
9      Set $\ell=\ell+1$
Algorithm 2 Derivative-Free Alternate Minimization
Clearly, the method revolves around the DF-Search subprocedure, described in Algorithm 3. In the latter, coordinate directions are explored according to classical derivative-free line search principles [20]: each direction (and its opposite) are polled for a given tentative stepsize, checking a sufficient decrease condition: if the step can be accepted, a further extrapolation is carried out as long as sufficient decrease is attained and the tentative stepsize at the subsequent iteration (for that direction) will be equal to the current actual stepsize; on the other hand, if sufficient decrease is not attained, the direction is discarded altogether for the iteration and the tentative stepsize is decreased for the next one.
1 Input: $g:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$, $x^{\ell},w^{1}\in\mathbb{R}^{n},$ $\tilde{\alpha}^{\ell}\in\mathbb{R}^{n}$, $\theta\in(0,1)$, $\gamma\in(0,1)$
2 Let $d_{1},\ldots,d_{n}=e_{1},\ldots,e_{n}$
3 for i=1,…,n do
4       if $g(x^{\ell},w^{i}+\tilde{\alpha}^{\ell}_{i}d_{i})\leq g(x^{\ell},w^{i})-\gamma(\tilde{\alpha}^{\ell}_{i})^{2}$ then
5            Set $\alpha_{i}=\max_{j=0,1,\ldots}\{\tilde{\alpha}^{\ell}_{i}\theta^{-j}\mid g(x^{\ell},w^{i}+\tilde{\alpha}^{\ell}_{i}\theta^{-j}d_{i})\leq g(x^{\ell},w^{i})-\gamma(\tilde{\alpha}^{\ell}_{i}\theta^{-j})^{2}\}$
6             Set $w^{i+1}=w^{i}+\alpha_{i}d_{i}$
7             Set $\tilde{\alpha}_{i}^{\ell+1}=\alpha_{i}$
8      else if $g(x^{\ell},w^{i}-\tilde{\alpha}^{\ell}_{i}d_{i})\leq g(x^{\ell},w^{i})-\gamma(\tilde{\alpha}^{\ell}_{i})^{2}$  then
9            Set $\alpha_{i}=\max_{j=0,1,\ldots}\{\tilde{\alpha}^{\ell}_{i}\theta^{-j}\mid g(x^{\ell},w^{i}-\tilde{\alpha}^{\ell}_{i}\theta^{-j}d_{i})\leq g(x^{\ell},w^{i})-\gamma(\tilde{\alpha}^{\ell}_{i}\theta^{-j})^{2}\}$
10             Set $w^{i+1}=w^{i}-\alpha_{i}d_{i}$
11             Set $\tilde{\alpha}_{i}^{\ell+1}=\alpha_{i}$
12      else
13             Set $\alpha_{i}=0$, $\tilde{\alpha}_{i}^{\ell+1}=\theta\tilde{\alpha}_{i}^{\ell}$
14
return $w^{n+1},\tilde{\alpha}^{\ell+1}$
Algorithm 3 Derivative-Free Search (DF-Search)
We can now turn to the analysis of the inner derivative-free optimization method.
3.1 Convergence analysis
We begin the analysis for Algorithms 2-3 stating the following preliminary results. Note that the property, stated for the $y$ block of variables, straightforwardly holds also for $z$.
Proposition 3.1.
Assume that the gradients of functions $f_{j}$, $j=1,\ldots,m$, are Lipschitz continuous over $\mathbb{R}^{n}$. Then, for any $\tau_{k}>0$, there exist a constants $c_{1},c_{2}>0$ such that for all $\ell$ we have
$\|\nabla_{y}P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})\|\leq(c_{1}+c_{2}\tau_{k})\max_{i=1,\ldots,n}\tilde{\alpha}_{yi}^{\ell+1}.$
Proof.
Let us consider a generic iteration $\ell$ and let us denote by $w^{\ell,i}$ the intermediate points visited in the DF-Search procedure called at step 2 in this iteration. Given any $i\in\{1,\ldots,n\}$, we can distinguish two cases:
  • (a)
    $w^{\ell,i+1}=w^{\ell,i}$;
  • (b)
    $w^{\ell,i+1}\neq w^{\ell,i}$.
Case (a). From the instructions of Algorithm 3, letting $g(x^{\ell},w)=P_{\tau_{k}}(x^{\ell},w,z^{\ell})$, we have
$g\left(x^{\ell},w^{\ell,i}\pm{{\tilde{\alpha}_{yi}^{\ell}}}d_{i}\right)>g(x^{\ell},w^{\ell,i})-\gamma\left({{\tilde{\alpha}_{yi}^{\ell}}}\right)^{2},$
i.e., since $\tilde{\alpha}_{yi}^{\ell+1}=\theta\tilde{\alpha}_{yi}^{\ell}$,
$g\left(x^{\ell},w^{\ell,i}\pm{{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}d_{i}\right)>g(x^{\ell},w^{\ell,i})-\gamma\left({{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}\right)^{2}.$
By applying the Mean Value Theorem in the above inequality in the case of direction $+d_{i}$, we have
$\nabla_{w}g(x^{\ell},u_{i}^{\ell})^{T}e_{i}\geq-\gamma\left({{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}\right),\quad u_{i}^{\ell}=w^{\ell,i}+t_{i}^{\ell}{{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}d_{i},\quad t_{i}^{\ell}\in(0,1)$ (14)
Recalling that functions $f_{j}$ are $L$-smooth and that $g(x^{\ell},w)=f_{1}(w)+\frac{\tau_{k}}{2}\|w-x^{\ell}\|^{2}+C$ where $C$ is some constant, we can deduce that $g$ also has also gradients $\nabla_{w}g(x^{\ell},w)$ that are Lipschitz continuous for some constant $L_{2}\leq L+\tau_{k}$ (independent of $x^{\ell}$). We can then write
$\nabla_{w}g(x^{\ell},u_{i}^{\ell})^{T}e_{i}$ $=\left(\nabla_{w}g(x^{\ell},u_{i}^{\ell})^{T}e_{i}-\nabla_{w}g(x^{\ell},y^{\ell})^{T}e_{i}+\nabla_{w}g(x^{\ell},y^{\ell})^{T}e_{i}\right)$
$\leq\nabla_{w_{i}}g(x^{\ell},y^{\ell})+\|\nabla_{w}g(x^{\ell},u_{i}^{\ell})-\nabla_{w}g(x^{\ell},y^{\ell})\|$
$\leq\nabla_{w_{i}}g(x^{\ell},y^{\ell})+(L+\tau_{k})\|u_{i}^{\ell}-y^{\ell}\|$
$\leq\nabla_{w_{i}}g(x^{\ell},y^{\ell})+(L+\tau_{k})\|w^{\ell,i}-y^{\ell}\|+(L+\tau_{k}){{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}$
$\leq\nabla_{w_{i}}g(x^{\ell},y^{\ell})+n(L+\tau_{k})\max_{h=1,\ldots,n}\{\tilde{\alpha}_{yh}^{\ell+1}\}+(L+\tau_{k}){{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}$
$\leq\nabla_{w_{i}}g(x^{\ell},y^{\ell})+\frac{(n+1)(L+\tau_{k})}{\theta}\max_{h=1,\ldots,n}\{\tilde{\alpha}_{yh}^{\ell+1}\},$
so that, by (14), it follows
$\nabla_{w_{i}}g(x^{\ell},y^{\ell})\geq-\frac{(n+1)(L+\tau_{k})+\gamma}{\theta}\max_{h=1,\ldots,n}\{\tilde{\alpha}_{yh}^{\ell+1}\}.$
By an analogous reasoning for the case of direction $-d_{i}$ we obtain
$\nabla_{w_{i}}g(x^{\ell},y^{\ell})\leq\frac{(n+1)(L+\tau_{k})+\gamma}{\theta}\max_{h=1,\ldots,n}\{\tilde{\alpha}_{yh}^{\ell+1}\}.$
We can therefore conclude that
$|\nabla_{w_{i}}g(x^{\ell},y^{\ell})|=|\nabla_{y_{i}}P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})|\leq\frac{(n+1)(L+\tau_{k})+\gamma}{\theta}\max_{h=1,\ldots,n}\{\tilde{\alpha}_{yh}^{\ell+1}\}.$ (15)
Case (b). Again, from the instructions of Algorithm 3 and assuming without loss of generality that sufficient decrease is obtained along $+d_{i}$, we have
$g(x^{\ell},w^{\ell,i}+\tilde{\alpha}_{yi}^{\ell+1}d_{i})$ $\leq g(x^{\ell},w^{\ell,i})-\gamma\left(\tilde{\alpha}_{yi}^{\ell+1}\right)^{2}$
$\leq g(x^{\ell},w^{\ell,i})+\gamma\left({{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}\right)^{2},$
and also
$g\left(x^{\ell},w^{\ell,i}+{{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}d_{i}\right)>g(x^{\ell},w^{\ell,i})-\gamma\left({{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}\right)^{2}.$
By the Mean Value Theorem, we get
$\nabla_{w}g(x^{\ell},u_{i}^{\ell})^{T}e_{i}\leq\gamma\left({{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}\right),\quad u_{i}^{\ell}=w^{\ell,i}+t_{i}^{\ell}{{\tilde{\alpha}_{yi}^{\ell+1}}}e_{i},\quad t_{i}^{\ell}\in(0,1),$
$\nabla_{w}g(x^{\ell},v_{i}^{\ell})^{T}e_{i}>-\gamma\left({{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}\right),\quad v_{i}^{\ell}=w^{\ell,i}+q_{i}^{\ell}{{\tilde{\alpha}_{yi}^{\ell+1}}\over{\theta}}e_{i},\quad q_{i}^{\ell}\in(0,1).$
Following a similar reasoning as in case (a), we obtain
$\nabla_{w_{i}}g(x^{\ell},y^{\ell})\leq\frac{(n+1)(L+\tau_{k})+\gamma}{\theta}\max_{h=1,\ldots,n}\{\tilde{\alpha}_{yh}^{\ell+1}\},$
$\nabla_{w_{i}}g(x^{\ell},y^{\ell})\geq-\frac{(n+1)(L+\tau_{k})+\gamma}{\theta}\max_{h=1,\ldots,n}\{\tilde{\alpha}_{yh}^{\ell+1}\},$
i.e.,
$|\nabla_{w_{i}}g(x^{\ell},y^{\ell})|=|\nabla_{y_{i}}P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})|\leq\frac{(n+1)(L+\tau_{k})+\gamma}{\theta}\max_{h=1,\ldots,n}\{\tilde{\alpha}_{yh}^{\ell+1}\}.$ (16)
Finally, from (15) and (16), we may conclude that
$\|\nabla_{y}P_{\tau_{k}}(y^{\ell},z^{\ell},x^{\ell})\|\leq\sqrt{n}\frac{(n+1)(L+\tau_{k})+\gamma}{\theta}\max_{h=1,\ldots,n}\{\tilde{\alpha}_{yi}^{\ell}\},$
and this concludes the proof.
We next prove a result concerning the sequences of stepsizes generated by using Algorithms 2-3. In particular, we show that if Algorithm 2 did not stop, the stepsizes would go to zero.
Proposition 3.2.
Assume that $\ell\to\infty$ in Algorithm 2. Then, for all $i=1,\ldots,n$, the following properties hold:
$\lim_{\ell\to\infty}\alpha_{yi}^{\ell}=0,\qquad\lim_{\ell\to\infty}\alpha_{zi}^{\ell}=0,$ (17)
$\lim_{\ell\to\infty}\tilde{\alpha}_{yi}^{\ell}=0,\qquad\lim_{\ell\to\infty}\tilde{\alpha}_{zi}^{\ell}=0.$ (18)
Proof.
The instructions of the algorithm imply that for all $\ell$ we have
$P_{\tau_{k}}(x^{\ell+1},y^{\ell+1},z^{\ell+1})\leq P_{\tau_{k}}(x^{\ell},y^{\ell+1},z^{\ell+1})\leq P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell}),$
and hence it follows that the sequence $\{P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})\}$ converges. Furthermore, by the instructions of Algorithm 3, the update of each individual variable from vectors $y$ and $z$ guarantees a sufficient decrease of $-\gamma{(\alpha_{i}^{\ell})}^{2}$ (this is trivially satisfied even when $\alpha_{i}^{\ell}=0$); we can therefore write
$P_{\tau_{k}}(x^{\ell},y^{\ell+1},z^{\ell+1})\leq P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})-\gamma\displaystyle{\sum_{i=1}^{n}}\left((\alpha_{yi}^{\ell})^{2}+(\alpha_{zi}^{\ell})^{2}\right).$
Taking into account the convergence of $\{P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})\}$, it follows that (17) holds.
We now prove (18); for the sake of simplicity we consider the sequence $\{\tilde{\alpha}^{\ell}_{y}\}$: the result for $\{\tilde{\alpha}^{\ell}_{z}\}$ can be obtained identically. For any $i\in\{1,\ldots,n\}$, we split the set of indices of iterates $\{0,1,\ldots\}$ into two subsets $T$ and $\bar{T}$ (for notational convenience we omit the dependence on $i$), where:
    $\ell\in T$ if and only if $\alpha_{yi}^{\ell}>0$;
    $\ell\in\bar{T}$ if and only if $\alpha_{yi}^{\ell}=0$.
For each $\ell\in T$ we have $\tilde{\alpha}_{yi}^{\ell+1}=\alpha_{yi}^{\ell}$, and hence, if $T$ is an infinite subset, from (17) it follows
$\lim_{\ell\in T,\ell\to\infty}\tilde{\alpha}_{yi}^{\ell+1}=0.$
For every $\ell\in\bar{T}$, let $m_{\ell}$ be the largest index such that $m_{\ell}<\ell$ with $m_{\ell}\in T$ (we assume $m_{\ell}=0$ if this index does not exist, i.e., $T$ is empty). We can write
$\tilde{\alpha}_{yi}^{\ell+1}=\theta^{\ell-m_{\ell}}\tilde{\alpha}_{yi}^{m_{\ell}+1}\leq\tilde{\alpha}_{yi}^{m_{\ell}+1}.$
For $\ell\in\bar{T}$ and $\ell\to\infty$ we have that either $m_{\ell}\to\infty$ (if $T$ is an infinite subset) or $\ell-m_{\ell}\to\infty$ (if $T$ is finite). In the former case, we have
$\lim_{\ell\in\bar{T},\ell\to\infty}\tilde{\alpha}_{yi}^{\ell+1}\leq\lim_{\ell\in\bar{T},\ell\to\infty}\tilde{\alpha}_{yi}^{m_{\ell}+1}=\lim_{\ell\in T,\ell\to\infty}\tilde{\alpha}_{yi}^{\ell+1}=0.$
In the latter case, letting $\bar{m}$ be the last index in $T$ and recalling that $\theta\in(0,1)$, we have
$\lim_{\ell\in\bar{T},\ell\to\infty}\tilde{\alpha}_{yi}^{\ell+1}=\lim_{\ell\in\bar{T},\ell\to\infty}\theta^{\ell-\bar{m}}\tilde{\alpha}_{yi}^{\bar{m}+1}=0.$
Putting everything together, we get (18). ∎
Finally, we show the convergence property w.r.t. the $x$ block of variables.
Proposition 3.3.
Assume that $\ell\to\infty$ in Algorithm 2. Then:
$\lim_{\ell\to\infty}\|x^{\ell}-\Pi_{X}[x^{\ell}-\nabla_{x}P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})]=0.$
Proof.
By the instructions of Algorithm 2, $\{P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})\}$ is a decreasing sequence, hence $\{(x^{\ell},y^{\ell},z^{\ell})\}\subseteq\mathcal{L}_{0}^{\tau_{k}}=\{(x,y,z)\mid P_{\tau_{k}}(x,y,z)\leq P_{\tau_{k}}(\hat{x}^{k-1},\hat{y}^{k-1},\hat{z}^{k-1})\}$. By Corollary 2.2, the level set $\mathcal{L}_{0}^{\tau_{k}}$ is compact, therefore the sequence $\{(x^{\ell},y^{\ell},z^{\ell})\}$ admits limit points. Let us now assume, by contradiction, that there exists an infinite subset ${T}$ such that
$\lim_{\ell\in{T},\ell\to\infty}(x^{\ell},y^{\ell},z^{\ell})=(\bar{x},\bar{y},\bar{z})$ (19)
with
$\|\bar{x}-\Pi_{X}[\bar{x}-\nabla_{x}P_{\tau_{k}}(\bar{x},\bar{y},\bar{z})]\|\geq\eta>0.$ (20)
From Proposition 3.2 it follows
$\lim_{\ell\in{T},\ell\to\infty}y^{\ell+1}$ $=\bar{y},$ (21)
$\lim_{\ell\in{T},\ell\to\infty}z^{\ell+1}$ $=\bar{z}.$
The instructions of Algorithm 2 imply
$P_{\tau_{k}}(x^{\ell+1},y^{\ell+1},z^{\ell+1})\leq P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})$
and
$P_{\tau_{k}}(x^{\ell+1},y^{\ell+1},z^{\ell+1})\leq P_{\tau_{k}}(x^{\ell}+\alpha_{\ell}d_{x}^{\ell},y^{\ell+1},z^{\ell+1})\leq P_{\tau_{k}}(x^{\ell},y^{\ell+1},z^{\ell+1}),$
where
$d_{x}^{\ell}=x^{\ell}-\Pi_{X}[x^{\ell}-\nabla_{x}P_{\tau_{k}}(x^{\ell},y^{\ell+1},z^{\ell+1})]$
and $\alpha_{\ell}$ is determined by the Armijo-type line search. From the properties of the Armijo-type line search ([12, Prop. 20.4]) we get
$\lim_{\ell\in{T},\ell\to\infty}\nabla_{x}P_{\tau_{k}}(x^{\ell},y^{\ell+1},z^{\ell+1})^{T}d_{x}^{\ell}=0.$
By the properties of the projected gradient direction [12, Eq. (20.15)], we also have
$\nabla_{x}P_{\tau_{k}}(x^{\ell},y^{\ell+1},z^{\ell+1})^{T}d_{x}^{\ell}\leq-\|d_{x}^{\ell}\|^{2}.$
We can therefore conclude that
$\lim_{\ell\in{T},\ell\to\infty}\|x^{\ell}-\Pi_{X}[x^{\ell}-\nabla_{x}P_{\tau_{k}}(x^{\ell},y^{\ell+1},z^{\ell+1})]\|=0.$
Then, recalling (19), (21) and the continuity properties of the projection mapping, we obtain
$\|\bar{x}-\Pi_{X}[\bar{x}-\nabla_{x}P_{\tau_{k}}(\bar{x},\bar{y},\bar{z})\|=0,$
and this contradicts (20). ∎
3.2 Inner stopping condition
While we proved the asymptotic convergence properties of Algorithm 2, we are in fact interested in suitably stopping it in finite time to use it within the sequential penalty scheme of Algorithm 1.
We are thus asked to devise a suitable stopping condition ensuring that the output $(x^{k},y^{k},z^{k})$ of Algorithm 2 satisfies conditions (5)-(7).
To this aim, we shall exploit the result from Proposition 3.1. We can indeed set as a termination criterion the combination of the following two properties:
$\max\left\{\max_{i=1,\ldots,n}\tilde{\alpha}_{yi}^{\ell+1},\max_{i=1,\ldots,n}\tilde{\alpha}_{zi}^{\ell+1}\right\}\leq$ $\frac{\xi_{k}}{\max\{\tau_{k},1\}},$ (22)
$\|x^{\ell}-\Pi_{X}[x^{\ell}-\nabla_{x}P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})]\|\leq$ $\;\xi_{k}.$
Note that the stopping condition should be checked after step 2 of Algorihtm 2 and, when satisfied, the point $(x^{\ell},y^{\ell},z^{\ell})$ should be given as output.
We first discuss the second condition. The gradient of $P_{\tau_{k}}$ w.r.t. $x$ variables is accessible, as it only concerns the quadratic penalty terms that we know in analytical form. Thus, the condition is easily checkable. Proposition 3.3 implies that the criterion is satisfied in a finite number of iterations.
As for the former condition, clearly it is readily checkable too, and by 3.2 it is also satisfied in a finite number of inner iterations. From Proposition 3.1 we also get
$\|\nabla_{y}P_{\tau_{k}}(x^{k},y^{k},z^{k})\|\leq(c_{1y}+c_{2}\tau_{k})\frac{\xi_{k}}{\max\{\tau_{k},1\}},$
$\|\nabla_{z}P_{\tau_{k}}(x^{k},y^{k},z^{k})\|\leq(c_{1z}+c_{2}\tau_{k})\frac{\xi_{k}}{\max\{\tau_{k},1\}}.$
If $\tau_{k}\geq 1$, we then get
$\|\nabla_{y}P_{\tau_{k}}(x^{k},y^{k},z^{k})\|\leq c_{1y}\frac{\xi_{k}}{\tau_{k}}+c_{2}\xi_{k}\leq(c_{1y}+c_{2})\xi_{k},$
$\|\nabla_{z}P_{\tau_{k}}(x^{k},y^{k},z^{k})\|\leq c_{1z}\frac{\xi_{k}}{\tau_{k}}+c_{2}\xi_{k}\leq(c_{1z}+c_{2})\xi_{k},$
whereas if $\tau_{k}<1$ we have
$\|\nabla_{y}P_{\tau_{k}}(x^{k},y^{k},z^{k})\|\leq c_{1y}{\xi_{k}}+c_{2}\tau_{k}\xi_{k}\leq(c_{1y}+c_{2})\xi_{k},$
$\|\nabla_{z}P_{\tau_{k}}(x^{k},y^{k},z^{k})\|\leq c_{1z}{\xi_{k}}+c_{2}\tau_{k}\xi_{k}\leq(c_{1z}+c_{2})\xi_{k},$
Thus, conditions (5)-(7) will be satisfied at each iteration $k$ and we may conclude that the convergence result of Proposition 2.4 holds.
4 The case of partially separable functions
The approach resulting from the combination of Sections 2 and 3 could in principle be employed to solve any problem of the form (1) without using the derivatives of $f$. However, as mentioned in the introduction of this manuscript, one of the computational motivations of the proposed strategy has to be found in the particular case of functions $f_{j}$ being coordinate partially separable. Formally, the problem has the form
$\min_{x\in X}\sum_{i=1}^{m}f_{j}(x_{S_{j}}),$ (23)
where $S_{j}\subseteq\{1,\ldots,n\}$ for all $j=1,\ldots,m$.
In this case, the reformulated problem for the penalty decomposition scheme is
$\min_{x\in X,y_{1},\ldots,y_{m}\in\mathbb{R}^{n}}$ $\sum_{i=1}^{m}f_{j}(y_{j})$
s.t. $x_{S_{j}}-y_{j}=0\quad\text{for all }j=1,\ldots,m.$
In this scenario, we have some advantages:
  • Each duplicate vector of variables is only made of few components.
  • As a consequence, the size of equality constraints to be penalized is heavily reduced.
  • Most importantly, in Algorithm 2, each DF-Search only concerns few variables, massively reducing the cost of derivative-free optimization steps; indeed, line-search base derivative-free approaches are know to scale poorly as the number of variables grow. The cost of an iteration of Algorithm 2 is now $\sum_{j=1}^{m}2|S_{j}|$ evaluations of individual functions $f_{j}$, which is equivalent to the cost of carrying out an iteration of the DF-Search procedure directly on the original problem;
  • In addition, the subproblems with respect to each block of variables $y_{j}$ are completely independent of each other; thus, optimization can naturally take advantage of parallel computation.
Note that the original problem should be harder to solve by (derivative-free) alternate minimization as functions $f_{j}$ are not separable. The variables of the overall problem are thus connected to each other by more complex functions than the penalty terms appearing in the subproblems.
5 Computational experiments
We report in this section an experimental evaluation of the proposed penalty decomposition derivative free (PDDF) method. We consider a set of problems of the form (23), where the functions $f_{j}$ are partially separable, by extending some test problems from CUTEst [11] as done in [21]. Each problem is obtained taking the summation of several copies of a base function, where a subset of variables is shared between functions, following different patterns. In ARWHEAD a variable is common to all the functions in the summation. In BDEXP, ENGVAL, MOREBV and TRIDIA variables are shared between consecutive functions, whereas in DIXMAANI, DIXMAANA, NZF1 and WOODS variables are shared following a more intricate rule. In BDQRTIC we find both variables shared between consecutive functions and a variable that is shared among them all. In BEALES, POWSING and ROSENBR instead, no variable is shared between the functions, thus each term of the summation can be optimized independently. We consider different dimensionalities $n$ for each problem, ranging from 10 to 6500. In Table 1 we report in detail the dimensionalities tested for each problem and the number of functions that compose the objective function. All problems considered in this experimental analysis are unconstrained.
Table 1: Details of the problems employed in the experimental analysis.
Problem $\mathbf{n}$ $\mathbf{m}$
$(n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})$
ARWHEAD 10, 50, 100, 500, 1000, 5000 $n-1$
BDEXP 10, 50, 100, 500, 1000, 5000 $n-2$
BDQRTIC 10, 50, 100, 500, 1000, 5000 $n-4$
BEALES 10, 50, 100, 500, 1000, 5000 $n/2$
BROYDN3D 10, 50, 100, 500, 1000, 5000 $n$
DIXMAANA 15, 51, 102, 501, 1002, 5001 $n$
DIXMAANI 15, 51, 102, 501, 1002, 5001 $n$
ENGVAL 10, 50, 100, 500, 1000, 5000 $n-1$
MOREBV 12, 52, 102, 502, 1002, 5002 $n$
NZF1 13, 39, 130, 650, 1300, 6500 $16n/13-$ 1
POWSING 20, 52, 100, 500, 1000, 5000 $n/4$
ROSENBR 10, 50, 100, 500, 1000, 5000 $n/2$
TRIDIA 10, 50, 100, 500, 1000, 5000 $n$
WOODS 20, 40, 200, 400, 2000, 4000 $23n/$ 2
The performance of the proposed PDDF method is evaluated in comparison to the classical line-search based derivative-free coordinate descent method (LS) proposed in [20], in its unconstrained variant: objective function sampling is carried out along coordinate directions, making extrapolation steps when a sufficient decrease of the objective function is obtained for the tentative stepsize. We refer the reader to [20] for further details.
The termination criterion for PDDF consists in the conditions (22) with $\xi_{k}=10^{-4}$ for all $k$, whereas for LS we employ $\max_{i=1,\dots,n}\alpha_{i}\leq 10^{-4}$. In case one algorithm does not meet the termination condition within 2 hours of wall-clock time or 10,000 outer cycle iterations, we stop the experiment. The initial stepsize is set to $1$ for both methods. As regards the line search, both methods use $\gamma=10^{-6}$ and $\theta=0.5$. For PDDF we set $\tau_{0}=f(x^{0})/(10^{2}m)$ and $\tau_{k+1}=\min\{1.05\tau_{k},f(x^{0})/m\}$, with $x^{0}\in\mathbb{R}^{n}$ being the initial solution of the problem considered. The upper bound on the penalty parameter is set for numerical stability reasons. The computational experiments here reported were implemented in Python3 and executed on a computer running Ubuntu 22.04 with an Intel Xeon Processor E5-2430 v2 6 cores 2.50 GHz and 32 GB RAM. The implementation of the proposed PDDF method is available at github.com/dadoPuccio/PD_derivative_free.
In order to offer a fair comparison between the two approaches, we count the number of function evaluations of each function $f_{j}$ that composes the objective function. By assuming that the computational burden for evaluating each $f_{j}$ is similar, the total of individual sub-function evaluations provides an unbiased metric to compare PDDF and LS. Table 2 reports the total count for PDDF and LS when their respective termination conditions are met. We observe that in most cases PDDF outperforms LS as it converges using fewer function evaluations. We also notice that LS scales very poorly with problem dimensionality: exploiting the problem structure, as in PDDF, does become crucial in this setting. Problem DIXMAANI results particularly challenging for LS, that is not capable to trigger the termination condition before the time limit is reached at any of the three largest dimensions we tested. In Table 3 we report the function values obtained by PDDF and LS when the termination condition is met. We observe that both algorithms are capable of obtaining high quality solutions, with the only exception being LS in the largest ROSENBR problem, as in that case the time limit is hit before the algorithm is capable to get a satisfactory solution.
Table 2: Total number of function evaluations for PDDF and LS when the termination condition is met. We use MAX_T and MAX_I to indicate cases where the maximum execution time or the maximum number of iterations was reached.
Problem $\mathbf{n}=n_{1}$ $\mathbf{n}=n_{2}$ $\mathbf{n}=n_{3}$ $\mathbf{n}=n_{4}$ $\mathbf{n}=n_{5}$ $\mathbf{n}=n_{6}$
PDDF LS PDDF LS PDDF LS PDDF LS PDDF LS PDDF LS
ARWHEAD 810 2709 4410 7.4e4 8910 3.0e5 4.5e4 7.5e6 9.0e4 3.0e7 4.5e5 7.5e8
BDEXP 7344 4424 4.2e4 1.4e5 8.6e4 5.6e5 4.4e5 1.4e7 8.7e5 5.7e7 4.4e6 1.4e9
BDQRTIC 4.1e4 3810 3.1e5 1.7e5 6.6e5 7.0e5 3.4e6 2.0e7 6.8e6 8.0e7 3.4e7 2.2e9
BEALES 600 2055 3000 5.1e4 6000 2.1e5 3.0e4 5.1e6 6.0e4 2.1e7 3.0e5 5.1e8
BROYDN3D 2.7e4 7200 1.8e5 1.6e5 4.2e5 6.4e5 2.1e6 1.6e7 4.1e6 6.4e7 2.1e7 1.6e9
DIXMAANA 3375 8640 1.1e4 1.0e5 2.3e4 4.0e5 1.2e5 9.6e6 2.4e5 3.8e7 1.3e6 9.6e8
DIXMAANI 6300 8625 2.1e4 1.0e5 5.2e4 4.0e5 3.0e5 MAX_T 1.7e6 MAX_T 1.3e7 MAX_T
ENGVAL 1.2e4 6174 6.5e4 1.8e5 1.3e5 7.3e5 6.5e5 1.8e7 1.3e6 7.3e7 6.5e6 1.8e9
MOREBV 1872 9.7e4 8372 3.8e5 1.6e4 9.2e5 7.2e4 9.6e6 1.4e5 3.8e7 6.7e5 7.5e8
NZF1 4875 3990 3.7e4 4.2e4 3.7e5 1.6e6 2.3e6 6.4e7 3.9e5 2.5e8 MAX_T MAX_T
POWSING 1160 3605 3016 2.4e4 5800 9.0e4 2.9e4 2.3e6 5.8e4 9.0e6 2.9e5 2.3e8
ROSENBR 4.3e4 7.3e4 2.2e5 1.8e6 4.3e5 7.3e6 2.2e6 1.8e8 4.3e6 7.3e8 2.2e7 MAX_T
TRIDIA 1.7e4 7350 1.3e5 2.5e5 4.3e5 1.1e6 1.1e7 2.7e7 4.7e7 1.1e8 MAX_T 2.8e9
WOODS 3690 2.3e4 7380 9.1e4 3.7e4 2.3e6 7.4e4 9.1e6 3.7e5 2.3e8 7.4e5 9.1e8
Table 3: Objective function values obtained with PDDF and LS when the termination condition is met.
Problem $\mathbf{n}=n_{1}$ $\mathbf{n}=n_{2}$ $\mathbf{n}=n_{3}$ $\mathbf{n}=n_{4}$ $\mathbf{n}=n_{5}$ $\mathbf{n}=n_{6}$
PDDF LS PDDF LS PDDF LS PDDF LS PDDF LS PDDF LS
ARWHEAD 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
BDEXP 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
BDQRTIC 11.9 11.9 106.0 106.0 223.7 223.7 1.2e3 1.2e3 2.3e3 2.3e3 1.2e4 1.2e4
BEALES 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
BROYDN3D 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DIXMAANA 15.0 15.0 51.0 51.0 102.0 102.0 501.0 501.0 1.0e3 1.0e3 5.0e3 5.0e3
DIXMAANI 15.0 15.0 51.0 51.0 102.0 102.0 501.0 501.0 1.0e3 1.0e3 5.0e3 5.0e3
ENGVAL 9.2 9.2 53.6 53.6 109.1 109.1 553.1 553.1 1.1e3 1.1e3 5.5e3 5.5e3
MOREBV 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
NZF1 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
POWSING 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
ROSENBR 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 74.2
TRIDIA 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.1 0.0
WOODS 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
We provide now the reader with some results concerning the wall-clock time that PDDF method employs to converge. Given the structure of the subproblems obtained with the penalty decomposition scheme, each duplicate vector of variables can be optimized independently of each other: therefore, a parallel implementation of the penalty decomposition derivative free method (PDDFP) can be fruitfully employed. This is particularly appealing for real-world black-box problems, where objective function evaluations are typically expensive to compute; hence, being able of evaluating multiple functions $f_{j}$ in parallel allows for considerable improvements to the overall computational time. The LS method, on the other hand, cannot exploit parallel computation; for a more challenging test, PDDFP is thus compared with a structure-aware implementation of the classical line-search based derivative-free coordinate descent method (SALS): by keeping track of the values of the individual functions $f_{j}$, it is possible to evaluate only a subset of the functions when a new point is considered, i.e., those impacted by changes to the currently considered variable. As regards the parallel implementation, we limit the maximum number of workers to 12. We simulate the high cost of function evaluations by introducing a 10ms pause at each evaluation. In Table 4 we report the outcome of this analysis. We observe that PDDFP consistently outperforms SALS when the problem dimension is large, as in that case the number of functions $f_{j}$ that constitute the objective is larger and the parallel implementation can be fruitfully employed.
Table 4: Wall-clock time required to reach the termination condition for the parallel version of PDDF and the structure-aware LS. Time is expressed in seconds.
Problem $\mathbf{n}=n_{1}$ $\mathbf{n}=n_{2}$ $\mathbf{n}=n_{3}$ $\mathbf{n}=n_{4}$
PDDFP SALS PDDFP SALS PDDFP SALS PDDFP SALS
ARWHEAD 1.1 5.6 3.0 30.4 5.1 61.5 21.8 311.1
BDEXP 11.2 13.7 21.6 83.8 48.2 171.4 204.1 873.6
BDQRTIC 79.7 19.3 159.0 174.7 310.7 352.4 1.5e3 2.0e3
BEALES 1.4 4.3 2.7 21.3 4.1 42.5 15.5 213.5
BROYDN3D 37.4 20.6 103.5 97.6 220.3 195.7 970.0 982.5
DIXMAANA 3.7 11.8 6.4 40.2 11.5 80.4 57.0 395.9
DIXMAANI 7.4 11.8 11.4 40.2 25.5 80.4 143.2 MAX_T
ENGVAL 18.0 12.7 40.9 73.9 72.0 149.3 309.3 754.6
MOREBV 2.0 14.8 5.3 72.3 8.4 138.3 33.7 495.3
NZF1 16.5 12.0 42.7 40.1 238.3 433.0 1.2e3 3.5e3
POWSING 2.6 7.2 2.7 18.8 5.1 36.2 15.8 181.5
ROSENBR 96.1 149.2 189.6 746.2 290.3 1.5e3 1.1e3 MAX_T
TRIDIA 23.0 14.4 75.6 100.1 222.5 215.8 5.1e3 1.1e3
WOODS 2.7 19.7 4.3 39.3 18.3 196.9 36.0 394.4
6 Conclusions
In this paper, we described and formally analyzed a penalty decomposition scheme for derivative-free optimization of coordinate partially separable finite-sum problems. The approach is proved to be well defined and possess global convergence guarantees under very mild standard assumptions. The results of preliminary computational experiments indicate that the proposed strategy shall be far superior than the blind employment of classical coordinate search approaches, especially as the size of the problems grows. Moreover, the approach is well-suited for parallelization; in scenarios where multi-processing is possible and the cost of computation is dominated by the evaluation of the black-box function, the proposed framework is shown to be superior also to a structure-informed smart implementation of the line-search based coordinate descent method.
The nice feedback obtained by the preliminary experiments opens up to interesting paths that could be explored in future research. First, working independently with each term in the objective should make the entire process less vulnerable to the presence of noise in a subset of the objectives. Moreover, the separability with respect to the sum terms is particularly nice in mixed setting where derivatives are available for some of the objectives [18]; indeed, gradient based optimization steps could be carried out w.r.t. blocks of variables associated with white-box terms; this would not be possible directly tackling the original formulation of the problem, where the same variables possibly appear in white and black box terms simultaneously. Finally, a strength point of penalty decomposition approaches in general is that they are able to handle additional constraints with a limited overhead (see, e.g., the discussion in [13]): indeed, additional constraints can be moved to the penalty function and handled naturally within the sequential penalty scheme.
Declarations
Funding
No funding was received for conducting this study.
Disclosure statement
The authors report there are no competing interests to declare.
Code availability statement
The implementation of the proposed PDDF method and its parallel variant is available at github.com/dadoPuccio/PD_derivative_free.
Notes on contributors
Francesco Cecere was born in Frosinone in 2000. He received his Bachelor’s degree in Mathematics in 2022 at Sapienza University of Rome. He then received his Master’s degree in Applied Mathematics in 2025 at the same university. He was an exchange student in 2023 at ENS Lyon.
Matteo Lapucci received his bachelor and master’s degree in Computer Engineering at the University of Florence in 2015 and 2018 respectively. He then received in 2022 his PhD degree in Smart Computing jointly from the Universities of Florence, Pisa and Siena. He is currently Assistant Professor at the Department of Information Engineering of the University of Florence. His main research interests include theory and algorithms for sparse, multi-objective and large scale nonlinear optimization.
Davide Pucci is a PhD student at the Global Optimization Laboratory of the University of Florence. He obtained both his Bachelor’s and Master’s degrees from the same university, in 2020 and 2023, respectively. His research focuses on optimization methods for machine learning and multi-objective optimization.
Marco Sciandrone is a full professor of Operations Research at Sapienza University of Rome. His research interests include operations research, nonlinear optimization, and machine learning. He has published more than 60 papers on international journals, and is coauthor of the book Introduction to Methods for Nonlinear Optimization, Springer, 2023.
ORCID
Matteo Lapucci: $0$-0002-2488-5486
Davide Pucci: $0$-0003-4424-9185
Marco Sciandrone: $0$-0002-5234-3174